OpenStudy (anonymous):

4x+5y=2 1/5x +y=-7/5

OpenStudy (anonymous):

solve one of the two equations for x or y. so suppose we solve for x in the first equation 4x + 5y = 2 4x = 2-5y x = (2/4) - (5/4)y x = 1/2 - (5/4)*y plug this value into the second equation for x. this will allow you to solve for y. then plug the value back in for y to solve for x.

OpenStudy (anonymous):

can you explain it again please